4(10pt). Find the general solution for the following equations. (a) y) – 4y(3) + 5y = 0.

(b) zy - xy - y=0. (Hint: Use substitution v = Inx.)
Answer #1

Solution :- y(4) _44 +5y=0 &2+54=0 y!4) Aurilliary Equation is n4-4 m3+5 m2=0 m m2-umts=0 m² (m²-4 m+5)=0 m=-(-4) Z √64=4x5Auxilliary equation is m²-2 m-l-o m=-(-2) I 18272-4 X1 X-1 2x at 4th 2 = 2118 = 2 + 252 F (1782) roots are distinct IJz if 52